ganeshie8 (ganeshie8):

A stone is dropped at t = 0. A second stone, with twice the mass of the first, is dropped from the same point at t = 100 ms. (a) How far below the release point is the center of mass of the two stones at t = 300 ms? (Neither stone has yet reached the ground.) (b) How fast is the center of mass of the two-stone system moving at that time?

OpenStudy (amistre64):

how do we calculate the center of mass?

ganeshie8 (ganeshie8):

\(\vec{r}_{cm} = \dfrac{\sum\limits_{i} m_i\vec{r}_i}{\sum\limits_{i}m_i}\)

ganeshie8 (ganeshie8):

differentiating gives \(\vec{v}_{cm} = \dfrac{\sum\limits_{i} m_i\vec{v}_i}{\sum\limits_{i}m_i}\)

OpenStudy (amistre64):

and where is the difficulty rising up at?

ganeshie8 (ganeshie8):

there is no difficulty as such my answer is not matching with the textbook, just wanted to take second opinion :) my answer is 5.7cm http://www.wolframalpha.com/input/?i=1%2F3%280.3%5E2%2B2*0.2%5E2%29 but textbook says 28cm

OpenStudy (amistre64):

is our distance from origin: d1 = gt^2/2 d2 = g(t-100)^2/2

ganeshie8 (ganeshie8):

yes, i used seconds, so i have below in my notes d1 = gt^2/2 d2 = g(t-0.1)^2/2

OpenStudy (amistre64):

then the distance between them is just d1-d2 right? center of mass to me means the point which would balance the 2 objects on a seesaw

ganeshie8 (ganeshie8):

Yes \[h_{cm} = \dfrac{d_1*(m) + d_2*(2m)}{m+2m} = \dfrac{0.3^2+2*0.2^2}{3}=5.7cm\]

ganeshie8 (ganeshie8):

Oh wait, I think I see the blunder... somehow I have cancelled g/2 wrongly

ganeshie8 (ganeshie8):

thanks @amistre64 that fixes the problem ! http://www.wolframalpha.com/input/?i=1%2F3%280.3%5E2%2B2*0.2%5E2%29*9.8%2F2

OpenStudy (amistre64):

yep, cant forget gravity :)

ganeshie8 (ganeshie8):

haha i had it in the equations, but i saw some imaginary "g/2" in the denominator and cancelled in that fraction... now both parts agree with the textbook answer !